OpenStudy (anonymous):

hey who can help with this: (a+5)^3−(b−5)^3/ (a+5)^4−(b−5)^4 can be written as A/B where A= B=

OpenStudy (anonymous):

anyone?

OpenStudy (tyteen4a03):

The equation is \(\frac{(a+5)^3−(b−5)^3} {(a+5)^4−(b−5)^4}\). Remember that \(\frac {a^x} {a^y}\) = a ^ (x - y).

OpenStudy (anonymous):

It's also useful to know that A^3 - B^3 = (A - B)(A^2 + AB + B^2) and the difference of squares.

OpenStudy (anonymous):

good point! Thanks guys!

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

still confused as to how to utilize this to solve it What would a^4-b^4 expand as?

OpenStudy (anonymous):

I utilized the rule a^z/a^y = a ^ (x - y) previously and came down with a-b+10- however, this cannot be the answer as the answer has to be in the form A/B :(

OpenStudy (anonymous):

It should expand as A^2 - B^2 = (A+B)(A-B), and A is in this case (a-5)^2 and B is (b-5)^2

OpenStudy (anonymous):

I apologize I switched the denominator and numerator (a+5)^4−(b−5)^4/ (a+5)^3−(b−5)^3 can be written as A/B where

OpenStudy (anonymous):

Doesn't matter, just apply the transformations we discussed.